# Give Me FP32 or Give Me Death? Challenges and Solutions for Reproducible Reasoning

### v1

@@ Appendix F Supplementary Results on Ablation Study
Tables 10 and 11 discuss the effect of different numbers of GPUs since, intuitively, the effect has a dependency on the GPU version, so we consider the situation on L40S and A100, separately. Table 10 reports the Avg_Std@top1_prob of LLM inference responses for the same question using 2 or 4 L40S GPUs under three different batch sizes (8, 16, and 32), while Table 11 presents the corresponding results using A100 GPUs. As shown in Table 10, increasing the number of GPUs from 2 to 4 generally leads to higher Avg_Std@top1_prob. This observation suggests that increasing the number of GPUs may introduce greater inference nondeterminism. However, this trend becomes less consistent in the A100 setting as shown in Table 11, where in many cases the 2 GPU results are even slightly higher than those of 4 GPUs. One reason behind it is that A100s do have higher instability in our experiments, which may have more influence beyond the number of GPUs (see Figure 6 (c)).
Table 10: Instability (Avg_Std@top1_prob (×$\times 10^{-4}$)) under different numbers of L40S GPUs
@@
Table 11: Instability (Avg_Std@top1_prob (×$\times 10^{-4}$)) under different numbers of A100 GPUs
@@
Tables 141516 and 17 present the Std@Acc for each ablation study. Unlike Avg_Std@top1_prob, it’s hard to observe clear trends or patterns from Std@Acc. We argue that Avg_Std@top1_prob serves as a more informative metric for ablation studies, as it directly reflects the numerical instability introduced by rounding error—i.e., fluctuations in the predicted probabilities of the same token. However, such fluctuations do not always lead to a token flip, since a flip only occurs when the variation at a specific position exceeds the original top-1 and top-2 probability gap, which is inherently a stochastic event. Moreover, even when a token flip happens—for example, replacing “the” with “that”—it may not necessarily cause the evaluation outcome to change from correct to incorrect or vice versa, and thus may not affect the final accuracy.
Table 12: Instability (Avg_Std@top1_prob (×$\times 10^{-4}$)) under different batch sizes
@@
Table 13: Instability (Avg_Std@top1_prob (×$\times 10^{-4}$)) under different GPU versions
@@
Appendix G Supplementary Results on LayerCast
We evaluate the LayerCast method proposed in Section 4 using the DeepSeek-R1-Distill-Qwen-7B model across five benchmarks: AIME’24, MATH500, LiveCodeBench-Easy, LiveCodeBench-Medium, and LiveCodeBench-Hard. These evaluations span different batch sizes (8, 16, 32), GPU counts (2 and 4), and numerical precision settings (BF16, FP32, and LayerCast). All LayerCast experiments follow the same experimental configuration as our main studies.
Table 18 reports both the accuracy and its standard deviation across different GPU count and batch size configurations. To facilitate comparison, results under FP32 and BF16 are also included. As expected, BF16 exhibits the most instability, with significantly higher standard deviations. LayerCast matches FP32 in terms of stability (often with zero or near-zero std) while preserving memory efficiency. These results corroborate our main findings in Section 4, demonstrating that LayerCast delivers reproducibility on par with FP32 while operating within a lower memory footprint.
Appendix H Supplementary Results on Random Sampling
@@
Tables 192120, and 22 summarize the Pass@1 accuracies and their standard deviation on DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, Qwen2.5-7B-Instruct, Llama-3.1-8B-Instruct, respectively. In most cases, FP32 yields the lowest standard deviations, reflecting greater stability across configurations. However, it is worth noticing that there are (4 out of 12) cases where FP16 results are more stable. This suggests that during random sampling, the two sources of randomness can interleave. We still urge researchers to sample more extensively to obtain more reproducible results.
Benchmark 2x A100 4x A100 Std@Acc
BS=8 BS=16 BS=32 BS=8 BS=16 BS=32
AIME’24
   LayerCast 0.4333 0.4333 0.4333 0.4333 0.4333 0.4333 0
   FP32 0.4333 0.4333 0.4333 0.4333 0.4333 0.4333 0
   BF16 0.4667 0.4667 0.3667 0.4333 0.5333 0.4667 0.0544
MATH500
   LayerCast 0.8680 0.8680 0.8680 0.8660 0.8660 0.8660 0.0011
   FP32 0.8680 0.8680 0.8680 0.8680 0.8660 0.8700 0.0013
   BF16 0.8700 0.8620 0.8580 0.8540 0.8860 0.8700 0.0114
LCB-Easy
   LayerCast 0.7308 0.7308 0.7363 0.7363 0.7418 0.7418 0.0049
   FP32 0.7363 0.7363 0.7363 0.7308 0.7363 0.7363 0.0022
   BF16 0.7198 0.7418 0.7253 0.7418 0.6978 0.7363 0.0169
LCB-Medium
   LayerCast 0.3058 0.3010 0.2961 0.2961 0.3010 0.3058 0.0043
   FP32 0.3058 0.3010 0.3058 0.3058 0.3058 0.3010 0.0025
   BF16 0.2816 0.2767 0.3204 0.2961 0.2864 0.2718 0.0176
LCB-Hard
   LayerCast 0.0488 0.0488 0.0569 0.0569 0.0488 0.0569 0.0044
   FP32 0.0488 0.0488 0.0488 0.0569 0.0569 0.0569 0.0044
   BF16 0.0569 0.0732 0.0569 0.0650 0.0569 0.0650 0.0066

### v2

@@ Appendix D Supplementary Results on Ablation Study
Tables 13 and 14 discuss the effect of different numbers of GPUs since, intuitively, the effect has a dependency on the GPU version, so we consider the situation on L40S and A100, separately. Table 13 reports the Avg_Std@top1_prob of LLM inference responses for the same question using 2 or 4 L40S GPUs under three different batch sizes (8, 16, and 32), while Table 14 presents the corresponding results using A100 GPUs. As shown in Table 13, increasing the number of GPUs from 2 to 4 generally leads to higher Avg_Std@top1_prob. This observation suggests that increasing the number of GPUs may introduce greater inference nondeterminism. However, this trend becomes less consistent in the A100 setting as shown in Table 14, where in many cases the 2 GPU results are even slightly higher than those of 4 GPUs. One reason behind it is that A100s do have higher instability in our experiments, which may have more influence beyond the number of GPUs (see Figure 6 (c)).
Table 13: Instability (Avg_Std@top1_prob ($\times 10^{-4}$)) under different numbers of L40S GPUs
@@
Table 14: Instability (Avg_Std@top1_prob ($\times 10^{-4}$)) under different numbers of A100 GPUs
@@
Tables 171819 and 20 present the Std@Acc for each ablation study. Unlike Avg_Std@top1_prob, it’s hard to observe clear trends or patterns from Std@Acc. We argue that Avg_Std@top1_prob serves as a more informative metric for ablation studies, as it directly reflects the numerical instability introduced by rounding error—i.e., fluctuations in the predicted probabilities of the same token. However, such fluctuations do not always lead to a token flip, since a flip only occurs when the variation at a specific position exceeds the original top-1 and top-2 probability gap, which is inherently a stochastic event. Moreover, even when a token flip happens—for example, replacing “the” with “that”—it may not necessarily cause the evaluation outcome to change from correct to incorrect or vice versa, and thus may not affect the final accuracy.
Table 15: Instability (Avg_Std@top1_prob ($\times 10^{-4}$)) under different batch sizes
@@
Table 16: Instability (Avg_Std@top1_prob ($\times 10^{-4}$)) under different GPU versions
@@
Appendix E Supplementary Results on Random Sampling
@@
Tables 212223, and 24 summarize the Pass@1 accuracies and their standard deviation on DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, Qwen2.5-7B-Instruct, Llama-3.1-8B-Instruct, respectively. In most cases, FP32 yields the lowest standard deviations, reflecting greater stability across configurations. However, it is worth noticing that there are (4 out of 12) cases where FP16 results are more stable. This suggests that during random sampling, the two sources of randomness can interleave. We still urge researchers to sample more extensively to obtain more reproducible results.
@@
Appendix F Supplementary Results on LayerCast
We evaluate the LayerCast method proposed in Section 4 across multiple models and benchmarks. These evaluations span different batch sizes (8, 16, 32), GPU counts (2 and 4), GPU types (A100 and L40S), and numerical precision settings (BF16, FP32, and LayerCast). All LayerCast experiments follow the same experimental configuration as our main studies.
It is important to note that LayerCast’s effectiveness is not merely empirical but has strong theoretical foundations. As we explain in Section 4, LayerCast performs all computations in FP32 precision, which fundamentally addresses the root cause of non-determinism we identified—the accumulation of rounding errors from lower-precision arithmetic. The only difference from pure FP32 inference is the just-in-time casting from BF16 storage to FP32 for computation, which is a deterministic operation that introduces no additional variance. The experiments we present here serve to demonstrate the practical applicability of this approach across different model architectures.
Table 25 reports the accuracy and standard deviation for DeepSeek-R1-Distill-Qwen-7B across five benchmarks. To facilitate comparison, results under FP32 and BF16 are also included. As expected, BF16 exhibits the most instability, with significantly higher standard deviations. LayerCast matches FP32 in terms of stability (often with zero or near-zero std) while preserving memory efficiency.
Benchmark 2x A100 4x A100 Std@Acc
BS=8 BS=16 BS=32 BS=8 BS=16 BS=32
AIME’24
   LayerCast 0.4333 0.4333 0.4333 0.4333 0.4333 0.4333 0
   FP32 0.4333 0.4333 0.4333 0.4333 0.4333 0.4333 0
   BF16 0.4667 0.4667 0.3667 0.4333 0.5333 0.4667 0.0544
MATH500
   LayerCast 0.8680 0.8680 0.8680 0.8660 0.8660 0.8660 0.0011
   FP32 0.8680 0.8680 0.8680 0.8680 0.8660 0.8700 0.0013
   BF16 0.8700 0.8620 0.8580 0.8540 0.8860 0.8700 0.0114
LCB-Easy
   LayerCast 0.7308 0.7308 0.7363 0.7363 0.7418 0.7418 0.0049
   FP32 0.7363 0.7363 0.7363 0.7308 0.7363 0.7363 0.0022
   BF16 0.7198 0.7418 0.7253 0.7418 0.6978 0.7363 0.0169
LCB-Medium
   LayerCast 0.3058 0.3010 0.2961 0.2961 0.3010 0.3058 0.0043
   FP32 0.3058 0.3010 0.3058 0.3058 0.3058 0.3010 0.0025
   BF16 0.2816 0.2767 0.3204 0.2961 0.2864 0.2718 0.0176
LCB-Hard
   LayerCast 0.0488 0.0488 0.0569 0.0569 0.0488 0.0569 0.0044
   FP32 0.0488 0.0488 0.0488 0.0569 0.0569 0.0569 0.0044
   BF16 0.0569 0.0732 0.0569 0.0650 0.0569 0.0650 0.0066
To further demonstrate LayerCast’s general effectiveness across different model architectures, Table 26 presents results for DeepSeek-R1-Distill-Llama-8B on the AIME’24 benchmark across all 12 runtime configurations. These results confirm that LayerCast consistently achieves perfect reproducibility with 0% Std@Acc (matching FP32) while maintaining memory efficiency benefits, following the same pattern observed with DeepSeek-R1-Distill-Qwen-7B. In contrast, BF16 shows significant variance with Std@Acc of 5.87% on A100 configurations and 3.44% on L40S configurations.
Table 26: Accuracy of DeepSeek-R1-Distill-Llama-8B on AIME’24 under different GPU counts, GPU types, precisions (including LayerCast), and batch sizes.
Precision 2x A100 4x A100
BS=8 BS=16 BS=32 BS=8 BS=16 BS=32
LayerCast 36.67% 36.67% 36.67% 36.67% 36.67% 36.67%
FP32 36.67% 36.67% 36.67% 36.67% 36.67% 36.67%
BF16 30.00% 30.00% 36.67% 23.33% 40.00% 30.00%
2x L40S 4x L40S
BS=8 BS=16 BS=32 BS=8 BS=16 BS=32
LayerCast 36.67% 36.67% 36.67% 36.67% 36.67% 36.67%
FP32 36.67% 36.67% 36.67% 36.67% 36.67% 36.67%
BF16 30.00% 36.67% 26.67% 33.33% 30.00% 30.00%
These results corroborate our main findings in Section 4, demonstrating that LayerCast delivers reproducibility on par with FP32 across different model architectures while operating within a lower memory footprint.